----- Original Message ----- From: http account <http@pii3.brahms.bnl.gov> To: <brahms-l@bnl.gov>; <olchansk@bnl.gov>; <videbaek@bnl.gov> Sent: Wednesday, June 21, 2000 7:03 AM Subject: Shift report 20000621 00:00-08:00 > > ------------------------------------------------------------ > Date: Wed Jun 21 07:55:30 2000, Shift: 00:00-08:00 > Supervisor: ykl, with: HH > Log book#2 pages 48 - 49 > > Shift Summary: > ------------------------------------------------------------ > 1.One hour of beaams at 64 MeV/c. Problems in injection, > extraction from AGS, controls, and synch between AGS & RHIC. > 2. Continued a run from the previous shift for an hour at 70 g. > which may have a few candidates for collisions. > 3. No detector problem.. > ------------------------------------------------------------ >
This archive was generated by hypermail 2b29 : Wed Jun 21 2000 - 11:16:03 EDT